# Nonsmooth Calabi–Yau structures for algebras and coalgebras

### v1

@@ 9.5. Simply connected spaces
We refine our Poincaré duality results when the spaces under consideration are simply connected. Note that a simply connected finitely dominated space is homotopy equivalent to a finite CW complex, since as in 9.24 its Wall finiteness obstruction necessarily vanishes. Recall from 8.10 the definition of a simply connected coalgebra.
@@
Since $X$ is finitely dominated, $C_{\bullet}\Omega X$ is smooth, and hence $C_{\bullet}X$ is a strongly proper coalgebra by 5.25. Moreover, since $X$ was simply connected, we may assume that $C_{\bullet}X$ is simply connected by 9.39. Recalling that $X$ is a $d$-Poincaré duality space if and only if $C_{\bullet}X$ is a $d$-Frobenius (equivalently symmetric) coalgebra (9.30), the equivalence of the first three statements follows from 8.13, since $C^{\bullet}X$ is the linear dual of $C_{\bullet}X$. The implication $(2)\implies(4)$ is 7.17. Since $X$ is homotopy equivalent to a finite CW complex, cellular cohomology gives us a finite dimensional algebra $A$ and a quasi-isomorphism $A\simeq C^{\bullet}X$. The implication $(4)\implies(2)$ now follows from 7.33. Finally, the equivalence of $(3)$ and $(5)$ is 7.31. ∎
@@
As in the proof of 9.40, $C_{\bullet}X$ is a proper coalgebra, and hence if $X$ is a Poincaré duality space then $C^{\bullet}X$ is a Frobenius algebra by 8.13. Moreover, a Frobenius algebra is Gorenstein by 7.17, which proves the first claim. The second claim follows from 9.40. ∎
@@
Let $X$ be a path connected finitely dominated topological space. Then for the statements of 9.40, statement (1) implies all the others. Statement (2) implies statement (4), and the converse is true if $C^{\bullet}X$ admits a finite dimensional model. Statements (3) and (5) are equivalent.

### v2

@@ 9.5. Simply connected spaces
We refine our Poincaré duality results when the spaces under consideration are simply connected. Note that a simply connected finitely dominated space is homotopy equivalent to a finite CW complex, since as in 9.26 its Wall finiteness obstruction necessarily vanishes. Recall from 8.11 the definition of a simply connected coalgebra.
@@
Since $X$ is finitely dominated, $C_{\bullet}\Omega X$ is smooth, and hence $C_{\bullet}X$ is a strongly proper coalgebra by 5.25. Moreover, since $X$ was simply connected, we may assume that $C_{\bullet}X$ is simply connected by 9.41. Recalling that $X$ is a $d$-Poincaré duality space if and only if $C_{\bullet}X$ is a $d$-Frobenius (equivalently symmetric) coalgebra (9.32), the equivalence of the first three statements follows from 8.14, since $C^{\bullet}X$ is the linear dual of $C_{\bullet}X$. The implication $(2)\implies(4)$ is 7.17. Since $X$ is homotopy equivalent to a finite CW complex, cellular cohomology gives us a finite dimensional algebra $A$ and a quasi-isomorphism $A\simeq C^{\bullet}X$. The implication $(4)\implies(2)$ now follows from 7.33. Finally, the equivalence of $(3)$ and $(5)$ is 7.31. ∎
@@
As in the proof of 9.42, $C_{\bullet}X$ is a proper coalgebra, and hence if $X$ is a Poincaré duality space then $C^{\bullet}X$ is a Frobenius algebra by 8.14. Moreover, a Frobenius algebra is Gorenstein by 7.17, which proves the first claim. The second claim follows from 9.42. ∎
@@
Let $X$ be a path connected finitely dominated topological space. Then for the statements of 9.42, statement (1) implies all the others. Statement (2) implies statement (4), and the converse is true if $C^{\bullet}X$ admits a finite dimensional model. Statements (3) and (5) are equivalent.